Find the divergence and curl of the f=3xyi-5zj+10xk
1
Expert's answer
2021-08-27T14:12:17-0400

"f = 3xyi -5zj+10xk
\\divf= \nabla .F
\\\text{where $\nabla = \frac{\partial}{\partial x}i$ + $\frac{\partial}{\partial y}j$ + $\frac{\partial}{\partial z}k$}
\\F= \frac{\partial}{\partial x}3xy - \frac{\partial}{\partial y}5z + \frac{\partial}{\partial z}10x
\\=3y
\\curl F = \nabla * F
\\ = \begin{vmatrix}i&j&k\\
\frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z}\\
3xy & -5z & 10x 
\end{vmatrix}
\\= 5i-10j-3xk"